OpenStudy (anonymous):

what is the range of f(x)=2x^2−20

jimthompson5910 (jim_thompson5910):

y = 2x^2 - 20 is the same as y = 2(x-0)^2 + (-20) so it's in the form y = a(x-h)^2 + k

jimthompson5910 (jim_thompson5910):

the vertex is (h,k) h = 0 k = -20 so the vertex is (0, -20) which means that the lowest point on this graph is (0, -20)

jimthompson5910 (jim_thompson5910):

any other point on y = 2x^2 - 20 will be above this lowest point and it stretches upwards forever to infinity

OpenStudy (anonymous):

so it would be from (-20,0) then

jimthompson5910 (jim_thompson5910):

[-20, infinity) actually

jimthompson5910 (jim_thompson5910):

basically, the range is any y value from -20 onward to +infinity

OpenStudy (anonymous):

than you

jimthompson5910 (jim_thompson5910):

yw